OpenStudy (anonymous):

Write the quadratic function in the form f(x) = a (x-h)^2 +k . Then, give the vertex of its graph. f(x) = -3x^2 +12x - 14

OpenStudy (anonymous):

ok so I know how to do the second part now I'm pretty sure but what is the first part referring to exactly? oh and this is my last one!!! @jim_thompson5910

OpenStudy (fibonaccichick666):

so make up some numbers for a,h, and k is all it is asking

OpenStudy (fibonaccichick666):

oh oh strike that last

OpenStudy (fibonaccichick666):

so it wants you to complete the square(factor) the function so it looks like the vertex form for a parabola

OpenStudy (anonymous):

so the second part I plug in to here to find the vertex right? \[(-\frac{ b }{ 2a },c-\frac{ b^2 }{ 4a })\]

OpenStudy (xapproachesinfinity):

complete the square in order to learn this better

OpenStudy (fibonaccichick666):

ah... well, if you complete the square, you know the vertex.

OpenStudy (fibonaccichick666):

because it is (h,k)

OpenStudy (xapproachesinfinity):

you find the the vertex by completing the square

OpenStudy (fibonaccichick666):

you don't need that(nor do I know if it's right, too many formulas to memorize)

OpenStudy (xapproachesinfinity):

i never memorized any formula in math, just learn where things come from

OpenStudy (xapproachesinfinity):

completing the square is where it came from so no harm in bring the it back to play the role :)

OpenStudy (anonymous):

I'm not sure what the square is exactly I've had like 6 questions like these so far but before this " see what you know" quiz i had never learned this stuff before so I'm still trying remember how this all works

OpenStudy (misty1212):

\[\huge y=a\overbrace{\color{red}{(x-h)^2}}^{\text{this is the square}}+k\]

OpenStudy (xapproachesinfinity):

\[f(x)=-3x^2+12x-4=-3(x^2-4x)-4=-3(x^2-4x+4)+12-4\] \[=-3(x-2)^2+8\]

OpenStudy (xapproachesinfinity):

that how you complete the square

OpenStudy (fibonaccichick666):

I walk away for 5 minutes and you're already just giving answers x, for real yo?

OpenStudy (xapproachesinfinity):

so vertex (2,8) the graph open down

OpenStudy (xapproachesinfinity):

hehe she said still have more question similar no harm to learn from this example seeing how it is worked completely

OpenStudy (fibonaccichick666):

This is her last one haha

OpenStudy (xapproachesinfinity):

hehe well i read in her replay she has 6 more or something at any rate, my intentions are clear :)

OpenStudy (anonymous):

this is a see what I know quiz lol so I'll be learning it more in the lesson I just didn't want to do terrible on it

OpenStudy (misty1212):

method 2 compute \[-\frac{b}{2a}=-\frac{12}{2\times -3}=2\] then go right to \[y=-3(x-2)^2+k\] to find \(k\) replace \(x\) by \(2\) in the original equation

OpenStudy (misty1212):

much snappier imho

OpenStudy (xapproachesinfinity):

oh bad bad please don't look for your quiz answers like that if i knew you wouldnt get any answer

OpenStudy (anonymous):

I have been working on this for about 10 hours straight though so my brain is also like fried at this point so I really!!! appriciate the help @misty1212 @xapproachesinfinity @FibonacciChick666

OpenStudy (fibonaccichick666):

lol it's ok, I'm gonna let you explain,also I'm mad cause this is a quiz(even if it's not a summative assessment). I'm tired. Oh old age, how the body doth wither and groan in the wind and sun day after day. For thou shalt know the pain of mountains, as they slowly crumble and fall upon death.

OpenStudy (anonymous):

I promise I've watched youtube videos I've read multiple explanations and looked up tons of formulas My intentions were and are not to just be given the answer out of laziness

OpenStudy (misty1212):

ten hours of math?

OpenStudy (misty1212):

go have a glass of wine

OpenStudy (anonymous):

yes I've been working on it alll day literally! lol

OpenStudy (xapproachesinfinity):

that's some quite a language lol

OpenStudy (misty1212):

next semester it will all be different...

OpenStudy (fibonaccichick666):

lol thanks x. And even with good intentions, it still gets a reprimand in my book. Finally, watch the khanacademy.org lessons, they will save you from 10 hour days. Take the assessment quiz, don't look anything up while taking it, and then watch from where it sits ya at. You'll then understand the material a heck of a lot better especially if you are a nontraditional student.

OpenStudy (anonymous):

khanacademy were some of the videos I watched. I honestly with all due respect do not see the problem with also having help and things explained through here and how it's different than watching a video or reading any other explanation some I needed help being guided through and to be pointed in the right direction or to have my answers checked to make sure they were either right or I was on the right track. Regardless I have had a long night so I appreciate the help and opinions have a good night